# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

### v1

@@ B.2 Efficiency-Equity Tradeoff
Per Capita Social Income Wealth Years Per Capita Social Income Wealth Years
GDP Welfare Gini Gini GDP Welfare Gini Gini
Free Market 1.37e+05 32.97 0.89 0.92 1.10 1.41e+05 334.79 0.90 0.93 1.00
Saez Tax 2.34e+12 73.82 0.21 0.38 300.00 6.35e+11 498.88 0.68 0.73 100.578
US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55 1.41e+05 351.17 0.89 0.93 1.00
AI Economist-BC 4.24e+12 97.24 0.54 0.52 299.55 N/A N/A N/A N/A N/A
AI Economist 1.26e+05 72.81 0.88 0.91 1.00 N/A N/A N/A N/A N/A
SMFG 1.01e+13 96.90 0.51 0.53 299.89 9.68e+12 975.15 0.52 0.51 300.00
B.3 Behavior Cloning Experiments
@@
This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), MF-MARL, and MADDPG network. The first two, as their network outputs are probability distributions, use negative log-likelihood loss (Figure 7 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 7 right). The loss convergence curve of behavior cloning is shown in Figure 7. It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP. The losses corresponding to MFRL and MADDPG can converge to below 0.1.
@@
Appendix C Saez tax
The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. The specific calculation method is as follows (Saez, 2001). The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates. The marginal tax rates denoted as $\tau(z)$, are expressed as a function of pretax income $z$, incorporating elements such as the income-dependent social welfare weight $G(z)$ and the local Pareto parameter $\alpha(z)$.
@@ Appendix C Saez tax
Appendix D Theoretical Proof
Proposition D.1.
The action value functions $Q_{i}$ and $\tilde{Q}_{t}$ are equivalent, i.e. $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})=\tilde{Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$, $\forall\mathbf{s}_{t}\in\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}$, $\forall\mathbf{a}_{t}\in\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}$, $\forall L_{t}^{f}\in\mathbb{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$.
@@ Appendix D Theoretical Proof
We commence by defining a partial order relation $\prec$ on $\mathcal{S}^{f}\times\mathcal{A}^{f}$. For any given vector $\mathbf{v}$, we construct a sorted counterpart $\tilde{\mathbf{v}}$, which maintains the same elements as $\mathbf{v}$ but orders them in ascending sequence based on $\prec$. For example, if $\mathbf{v}=\{(s_{1},a_{1}),(s_{2},a_{2}),(s_{3},a_{3})\}$, then $\tilde{\mathbf{v}}$ is arranged as $\{(s_{(1)},a_{(1)}),(s_{(2)},a_{(2)}),(s_{(3)},a_{(3)})\}$, where $(s_{(1)},a_{(1)})\prec(s_{(2)},a_{(2)})\prec(s_{(3)},a_{(3)})$). Exploiting the homogeneity of the follower agents, we can equate $Q(\mathbf{s}_{t},\mathbf{a}_{t})$ with $Q(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})$, where $\mathbf{v}$ encapsulates the joint state-action pairs of the followers, and $\tilde{\mathbf{v}}$ represents its sorted form. We further define $\mathcal{\mathbf{v}}$ as the collection of all such sorted vectors, and $\mathcal{L}$ as the ensemble of mean-field distributions. The empirical distribution of the followers’ state-action pairs, denoted as $L:\mathcal{V}\rightarrow\Delta(\mathcal{V})$, is expressed as $L^{f}_{t}(s,a)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\{(\tilde{s}_{i},\tilde{a}_{i})=(\tilde{s},\tilde{a})\}}$. The bijectiveness of the mapping $L$, ensured by the sorting process, leads to the conclusion that $Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}(s^{l}_{t},a^{l}_{t},\mathbf{v})=Q_{t}(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})=\tilde{Q}_{t}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$. ∎

### v2

@@ D.1 Training Curves for Dynamic SMFGs
Settings Algorithm GDP per Capita Social Welfare Income Gini Wealth Gini Years
N=100 without BC Rule-based+Random 1.41e+05 69.27 0.89 0.92 1.00
DDPG+Random 1.41e+05 70.91 0.88 0.92 1.00
MADDPG 4.93e+04 55.74 0.91 0.92 1.00
IDDPG 1.21e+05 83.09 0.88 0.92 1.00
MF-MARL 8.66e+07 82.02 0.82 0.83 75.75
SMFRL 9.59e+12 96.87 0.52 0.51 300.00
N=1000 without BC Rule-based+Random 1.42e+05 705.85 0.90 0.93 1.00
DDPG+Random 1.42e+05 654.37 0.90 0.93 1.00
MADDPG 3.37e+09 645.38 0.83 0.85 63.58
IDDPG 1.33e+05 874.53 0.89 0.92 1.00
MF-MARL 1.31e+05 834.93 0.88 0.92 1.50
SMFRL 1.10e+13 968.94 0.54 0.53 300.00
N=100 with BC Rule-based+Real 3.66e+11 79.23 0.52 0.53 217.45
DDPG+Real 2.03e+12 94.50 0.46 0.48 299.85
MADDPG 6.38e+12 93.89 0.57 0.58 268.53
IDDPG 7.41e+12 98.16 0.53 0.55 300.00
MF-MARL 5.44e+12 98.21 0.50 0.52 300.00
SMFRL 1.01e+13 96.90 0.51 0.53 299.89
N=1000 with BC Rule-based+Real 1.41e+05 334.79 0.90 0.93 1.00
DDPG+Real 4.92e+11 527.09 0.75 0.79 100.68
MADDPG 6.82e+12 954.88 0.56 0.62 278.50
IDDPG 2.79e+12 512.19 0.77 0.81 100.68
MF-MARL 1.13e+05 440.00 0.90 0.93 1.00
SMFRL 9.68e+12 975.15 0.52 0.51 300.00
Figure 4: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 & N=1000) and with behavior cloning (N=100-BC & N=1000-BC).
D.2 Training Curves for Various Tax Policies
D.3 Efficiency-Equity Tradeoff of Policies
D.4 Behavior Cloning Experiments
@@
We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5.3 to compare with SMFG Households. We use the statistical data from the 2022 Survey of Consumer Finances (SCF)  [^3] https://www.federalreserve.gov/econres/scfindex.htm. Details about the behavior cloning experiments are provided below. This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), MF-MARL, and MADDPG network. The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 7 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 7 right). The loss convergence curve of behavior cloning is shown in Figure 7. It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP. The losses corresponding to MFRL and MADDPG can converge to below 0.1.
@@
Appendix E Saez tax
The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. The specific calculation method is as follows [50]. The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates. The marginal tax rates denoted as $\tau(z)$, are expressed as a function of pretax income $z$, incorporating elements such as the income-dependent social welfare weight $G(z)$ and the local Pareto parameter $\alpha(z)$.
@@ Appendix E Saez tax
Appendix F Theoretical Proof
Proposition F.1.
The action value functions $Q_{i}$ and ${Q}_{t}$ are equivalent, i.e. $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})={Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$, $\forall\mathbf{s}_{t}\in\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}$, $\forall\mathbf{a}_{t}\in\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}$, $\forall L_{t}^{f}\in\mathbb{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$.
@@ Appendix F Theoretical Proof
We commence by defining a partial order relation $\prec$ on $\mathcal{S}^{f}\times\mathcal{A}^{f}$. For any given vector $\mathbf{v}$, we construct a sorted counterpart $\tilde{\mathbf{v}}$, which maintains the same elements as $\mathbf{v}$ but orders them in ascending sequence based on $\prec$. For example, if $\mathbf{v}=\{(s_{1},a_{1}),(s_{2},a_{2}),(s_{3},a_{3})\}$, then $\tilde{\mathbf{v}}$ is arranged as $\{(s_{(1)},a_{(1)}),(s_{(2)},a_{(2)}),(s_{(3)},a_{(3)})\}$, where $(s_{(1)},a_{(1)})\prec(s_{(2)},a_{(2)})\prec(s_{(3)},a_{(3)})$). Exploiting the homogeneity of the follower agents, we can equate $Q(\mathbf{s}_{t},\mathbf{a}_{t})$ with $Q(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})$, where $\mathbf{v}$ encapsulates the joint state-action pairs of the followers, and $\tilde{\mathbf{v}}$ represents its sorted form. We further define $\mathcal{\mathbf{v}}$ as the collection of all such sorted vectors, and $\mathcal{L}$ as the ensemble of mean-field distributions. The empirical distribution of the followers’ state-action pairs, denoted as $L:\mathcal{V}\rightarrow\Delta(\mathcal{V})$, is expressed as $L^{f}_{t}(s,a)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\{(\tilde{s}_{i},\tilde{a}_{i})=(\tilde{s},\tilde{a})\}}$. The bijectiveness of the mapping $L$, ensured by the sorting process, leads to the conclusion that $Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}(s^{l}_{t},a^{l}_{t},\mathbf{v})=Q_{t}(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})={Q}_{t}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$. ∎
Appendix G Economic Model Details
$S(W^{*}_{t})=D(W^{*}_{t}),\mathcal{L}_{t}=\sum^{N}_{i}e_{t}^{i}h_{t}^{i}.$
$K_{t+1}+B_{t+1}-A_{t+1}=\left(r_{t}+1\right)(K_{t}+B_{t}-A_{t})$
$Y_{t}=C_{t}+G_{t}+X_{t}$